OpenStudy (anonymous):

A fly buzzes past you at 6 m/s. The frequency of the buzz made by its wings is 162 Hz. Assume the speed of sound to be 334 m/s. What is the observed frequency of the hum as the bumblebee approaches you?

OpenStudy (mathteacher1729):

If I recall correctly, the (NON relativistic) formula for the Doppler shift is: \[\huge f_o/f_s = c/(c+v_s)\] \(f_0\)= observed frequency \(f_s\)= source frequency =162 Hz \(c\)=speed of sound = 334 m/s \(v_s\)=velocity of source (the fly) = 6 m/s So now you must solve for \(f_o\)

OpenStudy (anonymous):

328 Hz 168 Hz 156 Hz 165 Hz It would be 156 Hz right?

OpenStudy (mathteacher1729):

How did you obtain 156 Hz?

OpenStudy (anonymous):

f/162 = 334/ (334+6)

OpenStudy (mathteacher1729):

Ok. Then what? :)

OpenStudy (anonymous):

162 * 334 which is 54,108 then ..... (334+6) which is 340 . Then 54,108/340 which is 159.14

OpenStudy (mathteacher1729):

Yup! You got it. :)

OpenStudy (anonymous):

So it's 156Hz? That was easy! Thanks.